Find the length of the segment with endpoints (2,-1) (5,-6)
Answer by TimothyLamb(4379)

---
given:
(2,-1) (5,-6)
---
find distance between points
---
L = sqrt( dx^2 + dy^2 )
dx = 2 - 5 = -3
dy = -1 - (-6) = 5
---
L = sqrt( 9 + 25 )
L ~= 5.831
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php